OpenStudy (anonymous):

Can you add the square root of 6 and the square root of 24? Explain why or why not.

OpenStudy (raden):

yes, both numbers can be added see that sqrt(24) = sqrt(4) * sqrt(6) = 2 sqrt(6) so, sqrt(6) + sqrt(24) = sqrt(6) + 2 sqrt(6) = 3 sqrt(6) does that make sense ?